OpenStudy (anonymous):

5,-30,80,-1080, __, __ ? Help please i need the next two .

OpenStudy (mathstudent55):

The first step is to find if this is an arithmetic or a geometric sequence.

OpenStudy (mathstudent55):

To find which one it is, start by subtracting each number from the number to its right. If you get the same difference, it's an arithmetic sequence. What is -30 - 5? What is 80 - (-30)? If the differences are the same, it's an arithmetice sequence. If it isn't, then you need to check for a geometric sequence.

OpenStudy (mathstudent55):

To check for a geomteric sequence, instead of subtracting, you divide. Take each number and divide it by the previous number. If all quotients are the same, then it's a geometric sequence. The quotient between one number and the previous one is the ratio. Then just use that ratio and multiply it by the last number you are given to find the next one.

OpenStudy (anonymous):

Thanks, i think .

OpenStudy (mathstudent55):

BTW, you made a mistake in writing the problem. The third term is 180, not 80.

OpenStudy (mathstudent55):

5,-30,180,-1080 Check for an arithmetic sequence: -30 - 5 = -35 180 - (-30) = 180 + 30 = 210 -35 is clearly not equal to 210, so this is not an arithmetic sequence.

OpenStudy (mathstudent55):

5,-30,180,-1080 Check for a geometric sequence: -30/5 = -6 180 / (-30) = -6 -1080 / 180 = -6 All quotients are -6, so this is a geometric sequence. To find the next two terms, start by multiplying -1080 by -6. That the 5th term. Then multiply that result by -6 to get the 6th term.